# A Survey on LoRA of Large Language Models

### v1

@@ Abstract
Low-Rank Adaptation (LoRA), which updates the dense neural network layers with pluggable low-rank matrices, is one of the best performed parameter efficient fine-tuning paradigms. Furthermore, it has significant advantages in cross-task generalization and privacy-preserving. Hence, LoRA has gained much attention recently, and the number of related literature demonstrates exponential growth. It is necessary to conduct a comprehensive overview of the current progress on LoRA. This survey categorizes and reviews the progress from the perspectives of (1) downstream adaptation improving variants that improve LoRA’s performance on downstream tasks; (2) cross-task generalization methods that mix multiple LoRA plugins to achieve cross-task generalization; (3) efficiency-improving methods that boost the computation-efficiency of LoRA; (4) data privacy-preserving methods that use LoRA in federated learning; (5) application. Besides, this survey also discusses the future directions in this field.
@@ 1 Introduction
LoRA achieves parameter efficiency by updating the dense neural network layers of an LLM with pluggable low-rank matrices. These matrices (a.k.a, LoRA plugins) are independent of the LLM, which can be stored and reused in other related downstream tasks. Furthermore, these LoRA plugins can be combined to achieve cross-task generalization, which can facilitate multi-task learning, domain adaptation, and continual learning for LLMs.
As the LoRA plugins accumulate, the computation cost of managing LoRA plugins is increasing. Although LoRA is computation-efficient, the computational cost of managing a larger number of LoRA plugins is unignorable. It is necessary to further improve the computation efficiency of LoRA. The improvement can come from reducing the computation cost of single LoRA plugins and accelerating the scalable serving of multiple plugins. It can boost the application of LoRA in real-world use cases, such as Generative-as-a-Service (GaaS) cloud products.
In some cases, the training data are privately owned by multiple clients and cannot be centralized. To adapt LLMs with the distributed training data, we can adopt federated learning to protect the data privacy of each client. However, federated learning suffers expensive communication and computation costs. To reduce costs, LoRA is a natural choice. Its parameter-efficient nature helps to reduce the computation cost of each client and the communication cost of sharing parameters across clients. Furthermore, the pluggable feature of LoRA can help preserve the parameter privacy of each client in federated learning. Therefore, LoRA has a great potential for privacy-preserving.
In this survey, we give a comprehensive overview of the current progress on LoRA for methods (1) improving downstream adaption performance of LoRA; (2) mixing LoRA plugins to achieve cross-task generalization; (3) boosting the computation-efficiency of LoRA; (4) adopting LoRA in federated learning. Besides, the application of LoRA is briefly introduced. This taxonomy of LoRA-related methods is illustrated in Figure 1. This survey is expected to give comprehensive background knowledge, research trends and technical insights for LoRA.
The rest of this survey is organized as follows. Section 2 introduces the background knowledge of LoRA, and Section 3 introduces the LoRA’s variants that aim to improve the downstream adaptation performance. In Section 4, we review the LoRA mixture methods that mix LoRA plugins to achieve cross-task generalization. The LoRA-driven federated learning methods are introduced in Section 6. Section 7 reports the applications of LoRA. We conclude this survey and discuss the future directions in Section 8.
@@ 2 Low-Rank Adaptation (LoRA)
LoRAshear[84], Zhu et al.[85], leaf, text width=18em ] ] [ Parameter Sharing5.1.3) [ VeRA[86], VB-LoRA[87], leaf, text width=18em ] ] ] [ Parameter Quantization5.2), fill=blue!10 [ PTQ-based methods5.2.1) [ QLoRA[88], QA-LoRA[89], leaf, text width=18em ] ] [ QAT-base5.2.2) [ LoftQ[90], ApiQ[91], L4Q[92], leaf, text width=18em ] ] ] [ Parallel LoRA Computing Frameworks5.3), fill=blue!10 [ Parallel Fine-tuning5.3.1) [ ASPEN[93], leaf, text width=18em ] ] [ Parallel Inference5.3.2) [ Punica[94], S-LoRA[95], CARASERVE[96], leaf, text width=18em ] ] ] ] [ LoRA for Federate Learning6), fill=orange!10 [ Data Heterogeneity6.1), fill=orange!10 [ SLoRA [97], FeDeRA [98], FFA-LoRA [99], leaf, text width=20em ] ] [ Device Heterogeneity6.2), fill=orange!10 [ FedMS [100], FlexLoRA [101], HETLORA [102], leaf, text width=20em ] ] [ Model Heterogeneity6.3), fill=orange!10 [ pFedLoRA [103], leaf, text width=20em ] ] [ Parameter Privacy6.4), fill=orange!10 [ Huang et al. [104], PrivateLoRA [105], leaf, text width=20em ] ] ] [ Applications of LoRA7), fill=yellow!10 [ Language Tasks7.1), fill=yellow!10 [ Traditional NLP Task[106, 107, 108, 109, 110, 111, 112, 113, 114], Code Task[115, 116, 117, 118, 119, 120],
@@ 2.1 LoRA
LoRA is highly parameter efficient for it updates only a small subset of model parameters, which reduces the memory and computational requirements for fine-tuning without increasing inference latency [187]. Furthermore, The parameter efficiency can be further improved by extending from the low-rank matrix to low-rank tensor [188] or combining with the Kronecker decomposition [189, 190]. Except for parameter efficiency, LoRA is also pluggable for the LoRA parameters that can be separated from the model after training. The pluggable character of LoRA enables it to be shared and reused by multiple users  [191]. When we have LoRA plugins for multiple tasks, we can combine these plugins and expect a proper cross-task generalization performance [58]. Besides, the low-rank mechanism of LoRA is compatible with other parameter-efficient methods, such as adapter  [192, 193].
In practice, for a Transformer-based LLM, the dense layers typically consist of two types of weight matrices: the projection matrices in attention modules and feed-forward neural (FFN) modules. In the original study, LoRA is applied to the weight matrix of the attention layer. Subsequent work shows that using it in the FFN layers can further improve model performance [194].
@@ 2.2 Theoretical Analysis
As to the question that how LoRA can be more effective, Jang et al. [12] analyze the fine-tuning of LoRA within the neural tangent kernel (NTK) [198] framework, showing that employing a rank $r\gtrsim\sqrt{N}$ in LoRA helps to avoid spurious local minima and facilitates the discovery of low-rank solutions that exhibit good generalization. Besides, Zhu et al. [13] observe that the project-down matrix $A$ is utilized for extracting features from the input, while the project-up matrix $B$ employs these features to create the desired output. Based on this observation, they demonstrate that freezing the project-down matrix $A$ while tuning only the project-up matrix $B$ leads to better generalization compared to tuning both matrices, in addition to achieving a 2×2$\times$2 × reduction in parameters.
@@ 2.3 Efficiency in Practice
LoRA can significantly decrease the memory usage of fine-tuning an LLM, which can be divided into four parts: (1) Model Memory (Weight Memory): the memory required to store the model weights; (2) Activation Memory: the memory occupied by intermediate activations during forward propagation. It mainly depends on factors such as batch size and sequence length; (3) Gradient Memory: the memory required to store gradients during backpropagation. The gradients are only calculated for trainable parameters; (4) Optimization Memory: the memory used to store optimizer states. For example, the Adam optimizer stores the “first moment” and “second moment” of trainable parameters.
Literature [4] provides a comprehensive empirical comparison between full fine-tuning and LoRA fine-tuning on an LLaMA2-7B model with batch size 1, utilizing a single NVIDIA RTX4090 (24GB) GPU. According to this study, full fine-tuning requires approximately 60GB of memory, which exceeds the capacity of an RTX4090 GPU; by contrast, LoRA fine-tuning only needs about 23GB of memory. LoRA significantly reduces memory usage and makes fine-tuning LLaMA2-7B feasible on a single NVIDIA RTX4090 (24GB) GPU. Specifically, due to fewer trainable parameters, both optimization memory and gradient memory decrease significantly by approximately 25GB and 14GB respectively. On the other hand, while LoRA introduces additional “incremental parameters” resulting in slight increases in activation memory and weight memory (totaling about 2GB), this increase is negligible when considering the overall reduction in memory. Moreover, reducing memory brings an acceleration of forward propagation. LoRA is 1.9×1.9$\times$1.9 × times faster compared to full fine-tuning.
@@ 3 Downstream Adaptation Improving
Although LoRA can achieve proper adaptation performance on some downstream tasks, there is still a performance gap between LoRA and full fine-tuning on many downstream tasks, such as mathematical reasoning [200, 201, 202]. To fill this gap, many methods are proposed to further improve the downstream task adaption performance of LoRA. Typically, existing methods improve the downstream adaptation performance from the following perspectives: (1) breaking the low-rank bottleneck, refer to Figure 2 (c); (2) adaptively allocating the ranks of different LoRA modules, refer to Figure 2 (d); (3) optimizing the learning procedure of LoRA; (4) combining with other learning paradigms. In this section, we introduce these four types of methods respectively.
@@ 3.1.1 Stacking LoRAs along Fine-tuning
Matrix rank is subadditive, i.e., $rank(M_{1}+M_{2})\leq rank(M_{1})+rank(M_{2})$ for metrices $M_{1}$ and $M_{2}$ that have the same size. Based on the subadditivity, we can aggregate multiple LoRA modules together to increase the rank and break the low-rank bottleneck. Following this idea, ReLoRA [15] proposes a merge-and-reinit procedure for LoRA, which periodically merges the LoRA modules to the LLM and then reinitializes the LoRA modules during fine-tuning. It equals stacking multiple LoRA modules along with fine-tuning and can increase the rank of the overall updates. Similarly, COLA [22] proposes another merge-and-reinit method based on Frank-Wolfe algorithm [206]. However, MELoRA [23] points out that the merge-and-reinit procedure does not necessarily guarantee an increase in rank, because there can be overlap between the series of LoRA modules along fine-tuning. To solve this problem, MELoRA proposes to decompose the LoRA modules into smaller mini LoRAs and then parallelly stack these mini LoRAs, whose effectiveness in increasing the rank is theoretically verified.
@@ 3.1.3 Co-updating LLM and LoRA
The above two kinds of methods focus on improving the representation ability of LoRA itself. Different from them, Delta-LoRA [25] proposes to jointly update the LLM and LoRA modules, which directly updates the high-rank LLM and can gain better representations capable than updating LoRA independently. Delta-LoRA updates the LLM based on the difference between two LoRA modules of two consecutive iterations, which enables it to update the LLM without any extra memory.
@@ 3.3 Optimizing the Learning Procedure
In practice, LoRA converges more slowly than full fine-tuning. Moreover, it is also sensitive to hyperparameters and suffers from overfitting. These issues affect LoRA’s efficiency and hinder its downstream adaption performance. To address these issues, researchers have developed several approaches to optimize the learning procedure of LoRA, which can be categorized into the following three types: (1) Initialization Improvement; (2) Gradient Update Optimization; (3) Overfitting Mitigation.
@@ 4 Cross-task Generalization
[…]
@@ 4.1 Mixture with Manually Designed Weights
Early LoRA mixture methods attempt to linearly combine different LoRA plugins with manually designed weights. Some research demonstrates that we can achieve proper cross-task generalization ability by simply averaging the plugins or their related outputs [49, 50, 51]. Furthermore, several methods have been proposed to further improve the performance of the LoRA mixture via adopting manually designed weights. For example, ControlPE [52],  [53] and [54] set the weight factors as hyperparameters, and ControlPE uses hyperparameter search to determine the optimal combination of two LoRA plugins. Additionally, Token-level Adaptation [55] utilizes cosine similarity between the input feature and the adapter dataset center as weight factors, while BYOM [56] applies basic model fusion methods such as Task Arithmetic, Fisher-Merging, and RegMean.
@@ 4.2 Mixture with Learnt Weights
To learn the optimal mixture weights, several methods have been proposed at task level, instance level and token level to meet different needs. Task-level methods focus on enhancing task transferability, which can be either gradient-based, such as [57], or gradient-free, as seen in LoRAHub [58]. LoRAHub employs a black-box algorithm named CMA-ES [209] to optimize weight factors for LoRA plugins, simplifying the training process. Later, ComPEFT [59] and L-LoRA [60] use LoRAHub to mix quantized LoRA plugins, further improving computational efficiency.
Compared to task-level methods, instance-level and token-level methods can provide flexibility and precision for complex inputs. For multimodal instruction tuning, MixLoRA [61] dynamically chooses appropriate low-rank decomposition vectors based on the input instance, which are then integrated into LoRA matrices for training. To conduct protein mechanics analysis and design tasks, X-LoRA [62] develops a dynamic gating mechanism to assign weights for LoRA plugins at the token level and layer granularity. These approaches demonstrate better performance in specific tasks or application scenarios.

### v2

@@ Abstract
[…]
@@ 1 Introduction
LoRA achieves parameter efficiency by updating the dense neural network layers of an LLM with pluggable low-rank matrices. These matrices (a.k.a, LoRA plugins) are independent of the LLM, which can be stored and reused in other related downstream tasks. Furthermore, these LoRA plugins can be combined to achieve cross-task generalization, which can facilitate multi-task learning, domain adaptation, and continual learning.
As the LoRA modules accumulate, the computation cost of managing LoRA modules is increasing. Although LoRA is computation-efficient, the computational cost of managing a larger number of LoRA modules is unignorable. It is necessary to further improve the computation efficiency of LoRA. The improvement can come from reducing the computation cost of single LoRA modules and accelerating the scalable serving of multiple modules. It can boost the application of LoRA in real-world use cases, such as Generative-as-a-Service (GaaS) cloud products.
In some cases, the training data are privately owned by multiple clients and cannot be centralized. To adapt LLMs with the distributed training data, we can adopt federated learning to protect the data privacy of each client. However, federated learning suffers expensive communication and computation costs. To reduce costs, LoRA is a natural choice. Its parameter-efficient nature helps to reduce the computation cost of each client and the communication cost of sharing parameters across clients. Furthermore, the pluggable feature of LoRA, which supports the localization or encryption of personalized parameters, enhances privacy protection within federated learning. Therefore, LoRA has a great potential for privacy-preserving.
While some previous surveys have mentioned LoRA [5, 10, 11], they mainly focus on PEFT and only introduce a small number of LoRA-related works, lacking systematic treatment and comprehensive overview on LoRA and its variants. In this survey, we give a comprehensive overview of the current progress on LoRA for methods (1) improving downstream adaption performance of LoRA; (2) mixing LoRA modules to achieve cross-task generalization; (3) boosting the computation-efficiency of LoRA; (4) adopting LoRA in federated learning. Besides, the application of LoRA is briefly introduced. This taxonomy of LoRA-related methods is illustrated in Figure 1. This survey is expected to give comprehensive background knowledge, research trends and technical insights for LoRA.
The rest of this survey is organized as follows. Section 2 introduces the background knowledge of LoRA, and Section 3 introduces the LoRA’s variants that aim to improve the downstream adaptation performance. In Section 4, we review the LoRA mixture methods that mix LoRA modules to achieve cross-task generalization. The LoRA-driven federated learning methods are introduced in Section 6. Section 7 reports the applications of LoRA. We conclude this survey and discuss the future directions in Section 8.
@@ 2 Low-Rank Adaptation (LoRA)
LoRAshear[86], Zhu et al.[87], leaf, text width=18em ] ] [ Parameter Sharing5.1.3) [ VeRA[88], VB-LoRA[89], FourierFT [90], leaf, text width=18em ] ] ] [ Parameter Quantization5.2), fill=blue!10 [ PTQ-based methods5.2.1) [ QLoRA[91], QA-LoRA[92], leaf, text width=18em ] ] [ QAT-base5.2.2) [ LoftQ[93], ApiQ[94], L4Q[95], leaf, text width=18em ] ] ] [ Parallel LoRA Computing Frameworks5.3), fill=blue!10 [ Parallel Fine-tuning5.3.1) [ ASPEN[96], leaf, text width=18em ] ] [ Parallel Inference5.3.2) [ Punica[97], S-LoRA[98], CARASERVE[99], leaf, text width=18em ] ] ] ] [ LoRA for Federate Learning6), fill=orange!10 [ Data Heterogeneity6.1), fill=orange!10 [ SLoRA [100], FeDeRA [101], FFA-LoRA [102], leaf, text width=20em ] ] [ Device Heterogeneity6.2), fill=orange!10 [ FedMS [103], FlexLoRA [104], HETLORA [105], leaf, text width=20em ] ] [ Model Heterogeneity6.3), fill=orange!10 [ pFedLoRA [106], leaf, text width=20em ] ] [ Parameter Privacy6.4), fill=orange!10 [ Huang et al. [107], PrivateLoRA [108], leaf, text width=20em ] ] ] [ Applications of LoRA7), fill=yellow!10 [ Language Tasks7.1), fill=yellow!10 [ Traditional NLP Task[109, 110, 111, 112, 113, 114, 115, 116, 117], Code Task[118, 119, 120, 121, 122, 123],
@@ 2.1 LoRA
LoRA is highly parameter efficient for it updates only a small subset of model parameters, which reduces the memory and computational requirements for fine-tuning without increasing inference latency [190]. Furthermore, The parameter efficiency can be further improved by extending from the low-rank matrix to low-rank tensor [191] or combining with the Kronecker decomposition [192, 193]. Except for parameter efficiency, LoRA is also pluggable for the LoRA parameters that can be separated from the model after training. The pluggable character of LoRA enables it to be shared and reused by multiple users  [194]. When we have LoRA modules for multiple tasks, we can combine these modules and expect a proper cross-task generalization performance [60]. Besides, the low-rank mechanism of LoRA is compatible with other parameter-efficient methods, such as adapter  [195, 196].
Besides, LoRA can achieve proper downstream adaptation performance on various downstream tasks. For example, on MMLU [197] benchmark, comparing with full fine-tuning, fine-tuning with LoRA can achieve comparable or even better performance across 57 tasks [4].
In practice, for a Transformer-based LLM, the dense layers typically consist of two types of weight matrices: the projection matrices in attention modules and feed-forward neural (FFN) modules. The experiments mentioned above are conducted based on the original LoRA settings, applying it to the query and value weight matrices in the attention modules. It is worth mentioning that subsequent work shows that applying it to the FFN layers can further improve model performance [198].
@@ 2.2 Theoretical Analysis
As to the question that how LoRA can be more effective, Jang et al. [14] analyze the fine-tuning of LoRA within the neural tangent kernel (NTK) [202] framework when $N$ data points are available. They demonstrate that employing a rank $r\gtrsim\sqrt{N}$ in LoRA helps to avoid spurious local minima and facilitates the discovery of low-rank solutions that exhibit good generalization. Besides, Zhu et al. [15] observe that the project-down matrix $A$ is utilized for extracting features from the input, while the project-up matrix $B$ employs these features to create the desired output. Based on this observation, they demonstrate that freezing the project-down matrix $A$ while tuning only the project-up matrix $B$ leads to better generalization compared to tuning both matrices, in addition to achieving a 2×2$\times$2 × reduction in parameters.
@@ 2.3 Efficiency in Practice
LoRA can significantly decrease the memory usage of fine-tuning an LLM, which can be divided into four parts: (1) Model Memory: the memory required to store the model weights; (2) Activation Memory: the memory occupied by intermediate activations during forward propagation. It mainly depends on factors such as batch size and sequence length; (3) Gradient Memory: the memory required to store gradients during backpropagation. The gradients are only calculated for trainable parameters; (4) Optimization Memory: the memory used to store optimizer states. For example, the Adam optimizer stores the “first moment” and “second moment” of trainable parameters.
Pan et al. [4] provides a comprehensive empirical comparison between full fine-tuning and LoRA fine-tuning on an LLaMA2-7B model with batch size 1, utilizing a single NVIDIA RTX4090 (24GB) GPU. According to this study, full fine-tuning requires approximately 60GB of memory, which exceeds the capacity of an RTX4090 GPU; by contrast, LoRA fine-tuning only needs about 23GB of memory. LoRA significantly reduces memory usage and makes fine-tuning LLaMA2-7B feasible on a single NVIDIA RTX4090 (24GB) GPU. Specifically, due to fewer trainable parameters, both optimization memory and gradient memory decrease significantly by approximately 25GB and 14GB respectively. On the other hand, while LoRA introduces additional “incremental parameters” resulting in slight increases in activation memory and weight memory (totaling about 2GB), this increase is negligible when considering the overall reduction in memory. Moreover, reducing memory brings an acceleration of forward propagation. LoRA is 1.9×1.9$\times$1.9 × times faster compared to full fine-tuning.
@@ 3 Downstream Adaptation Improving
Although LoRA can achieve proper adaptation performance on some downstream tasks, there is still a performance gap between LoRA and full fine-tuning on many downstream tasks, such as mathematical reasoning [204, 205, 206]. To fill this gap, many methods are proposed to further improve the downstream adaptation performance of LoRA. Typically, existing methods improve the downstream adaptation performance from the following perspectives: (1) breaking the low-rank bottleneck, refer to Figure 2 (c); (2) adaptively allocating the ranks of different LoRA modules, refer to Figure 2 (d); (3) optimizing the learning procedure of LoRA; (4) combining with other learning paradigms. In this section, we introduce these four types of methods respectively.
@@ 3.1.1 Stacking LoRAs along Fine-tuning
Matrix rank is subadditive, i.e., $rank(M_{1}+M_{2})\leq rank(M_{1})+rank(M_{2})$ for matrices $M_{1}$ and $M_{2}$ that have the same size. Based on the subadditivity, we can aggregate multiple LoRA modules together to increase the rank and break the low-rank bottleneck. Following this idea, ReLoRA [17] proposes a merge-and-reinit procedure for LoRA, which periodically merges the LoRA modules to the LLM and then reinitializes the LoRA modules during fine-tuning. It equals stacking multiple LoRA modules along with fine-tuning and can increase the rank of the overall updates. Similarly, COLA [24] proposes another merge-and-reinit method based on Frank-Wolfe algorithm [210]. However, MELoRA [25] points out that the merge-and-reinit procedure does not necessarily guarantee an increase in rank, because there can be overlap between the series of LoRA modules along fine-tuning. To solve this problem, MELoRA proposes to decompose the LoRA modules into smaller mini LoRAs and then parallelly stack these mini LoRAs, whose effectiveness in increasing the rank is theoretically verified.
@@ 3.1.3 Co-updating LLM and LoRA
The above two kinds of methods focus on improving the representation ability of LoRA itself. Different from them, Delta-LoRA [27] proposes to jointly update the LLM and LoRA modules, which directly updates the high-rank LLM and can gain better representation capable than updating LoRA independently. It updates the LLM based on the difference between two LoRA modules of two consecutive iterations, which enables it to update the LLM without any extra memory.
@@ 3.3 Optimizing the Learning Procedure
In practice, LoRA converges more slowly than full fine-tuning. Moreover, it is also sensitive to hyperparameters and suffers from overfitting. These issues affect LoRA’s efficiency and hinder its downstream adaptation performance. To address these issues, researchers have developed several approaches to optimize the learning procedure of LoRA, which can be categorized into the following three types: (1) Initialization Improvement; (2) Gradient Update Optimization; (3) Overfitting Mitigation.
@@ 3.4 Combining with other Learning Paradigms
At last, to illustrate the performance difference between LoRA and some of its variants, we report their performance for RoBERTa-base [213] model on the GLUE benchmark [214] in Table 1. These results are derived from previous studies [9, 90, 32, 16, 45].
Table 1: Performance of LoRA and its variants for RoBERTa-base model on the GLUE benchmark. We report Matthew’s correlation for CoLA, Pearson correlation for STS-B, and accuracy for the other datasets. The results are reported according to the results reported in literature [9, 90, 32, 45, 89].
Method # Params SST-2 MPRC CoLA QNLI RTE STS-B
Tied-LoRA [215] 0.043M 94.4 88.5 61.9 92.0 76.2 89.8
AutoLoRA [32] 0.3M 94.9 89.4 61.3 92.9 77.0 90.8
DyLoRA [35] 0.3M 94.3 89.5 61.1 92.2 78.7 91.1
AdaLoRA [28] 0.3M 94.5 88.7 62.0 93.1 81.0 90.5
FourierFT [90] 0.024M 94.2 90.0 63.8 92.2 79.1 90.8
VeRA [88] 0.043M 94.6 89.5 65.6 91.8 78.7 90.7
Full Fine-tuning [9] 125M 94.8 90.2 63.6 92.8 78.7 91.2
LoRA [9] 0.3M 95.1 89.7 63.4 93.3 78.4 91.5
VB-LoRA [89] 0.023M 94.4 89.5 63.3 92.2 82.3 90.8
BiLoRA [45] 0.3M 95.1 91.7 64.8 93.3 87.2 91.7
@@ 4 Cross-task Generalization
LoRA’s pluggable nature enables users to accumulate LoRA plugins for different tasks. For example, on Hugging Face platform, there are more than 20,000 LoRA plugins compatible with various LLMs for different tasks. These accumulated LoRA plugins can not only be utilized independently but also be mixed to achieve cross-task generalization[60]. Mixing multiple LoRA plugins together, namely LoRA mixture, has been widely applied in areas requiring cross-task generalization, such as multi-task learning, domain adaptation, and continual learning. Existing LoRA mixture methods can be categorized into (1) mixture with manually designed weights; (2) mixture with learnt weights; (3) mixture of LoRA experts. This section introduces each category of methods respectively, as shown in Fig. 3.
@@ 4.1 Mixture with Manually Designed Weights
Early LoRA mixture methods attempt to linearly combine different LoRA modules with manually designed weights. Some research demonstrates that we can achieve proper cross-task generalization ability by simply averaging LoRA modules or their related outputs [51, 52, 53]. Furthermore, several methods have been proposed to further improve the performance of the LoRA mixture via adopting manually designed weights. For example, ControlPE [54],  [55] and [56] set the weight factors as hyperparameters, and ControlPE uses hyperparameter search to determine the optimal combination of two LoRA modules. Additionally, Token-level Adaptation [57] utilizes cosine similarity between the input feature and the adapter dataset center as weight factors, while BYOM [58] applies basic model fusion methods such as Task Arithmetic, Fisher-Merging, and RegMean.
@@ 4.2 Mixture with Learnt Weights
To learn the optimal mixture weights, several methods have been proposed at task level, instance level and token level to meet different needs. Task-level methods focus on enhancing task transferability, which can be either gradient-based, such as [59], or gradient-free, as seen in LoRAHub [60]. LoRAHub employs a black-box algorithm named CMA-ES [216] to optimize weight factors for LoRA modules, simplifying the training process. Later, ComPEFT [61] and L-LoRA [62] use LoRAHub to mix quantized LoRA modules, further improving computational efficiency.
Compared to task-level methods, instance-level and token-level methods can provide flexibility and precision for complex inputs. For multimodal instruction tuning, MixLoRA [63] dynamically chooses appropriate low-rank decomposition vectors based on the input instance, which are then integrated into LoRA matrices for training. To conduct protein mechanics analysis and design tasks, X-LoRA [64] develops a dynamic gating mechanism to assign weights for LoRA modules at the token level and layer granularity. These approaches demonstrate better performance in specific tasks or application scenarios.